# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

### v3

@@ Abstract
Analog electronic circuits are at the core of an important category of musical devices, which includes a broad range of sound synthesizers and audio effects. The nonlinear features of their passive and active electronic components give analog musical devices a distinctive timbre and sound quality, making them highly desirable. The development of software that simulates analog musical devices, known as virtual analog modeling, is a significant sub-field in audio signal processing. Artificial neural networks are a promising technique for virtual analog modeling. They have rapidly gained popularity for the emulation of analog audio effects circuits, particularly recurrent networks. While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response. Although hybrid solutions can offer advantages, black-box approaches can still be advantageous in some contexts. In this article, we explore the application of recent machine learning advancements for virtual analog modeling. In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks. These have shown promising ability in sequence-to-sequence modeling tasks, showing a notable improvement in signal history encoding. Our comparative study uses these black-box neural modeling techniques with a variety of audio effects. We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development. Our metrics aim to assess the models’ ability to accurately replicate energy envelopes and frequency contents, with a particular focus on transients in the audio signal. To incorporate control parameters into the models, we employ the Feature-wise Linear Modulation method. Long Short-Term Memory networks exhibit better accuracy in emulating distortions and equalizers, while the State-Space model, followed by Long Short-Term Memory networks when integrated in an encoder-decoder structure, and Linear Recurrent Unit outperforms others in emulating saturation and compression. When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history. The Long Short-Term Memory networks present more tendency to introduce audio artifacts.
keywords:
Nonlinear Modeling, Virtual Analog, Audio Effect, Machine Learning
@@ 1 Introduction
The emulation of analog musical devices has become incredibly pervasive in digital commercial products. The distinct sound characteristics of both contemporary and old analog audio devices have captivated producers and musicians. nalog effect exists as standalone units as well as integrated with sound synthesizers. There are different types of popular effects such as filters, distortion, overdrive, compressor/limiter, as well as delay-based effects like chorus, flanger, phaser, and reverb. Analog circuits found in audio effects depend on intricately designed circuits that combine discrete passive and active electronic components to generate and modify electrical signals. Key components, such as operational amplifiers and diodes, are utilized in a variety of analog effects. These components present nonlinear behaviors that contribute to the unique sound of these devices. Accurate modeling of such nonlinearities is essential to replicate the behavior of analog audio devices. Consequently, developing digital audio effects capable of simulating analog counterparts has become a crucial and valuable area of study within digital audio signal processing, often referred to as Virtual Analog (VA) modeling.
@@
This article presents a comparative study that applies recent advancements in machine learning to audio effects modeling. Machine learning has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks. Although hybrid solutions (referred to as gray-box approaches) have also been proposed, which offer their advantages  [1], black-box approaches can still be appealing in some contexts. Gray-box approaches can incorporate physics knowledge  [2] or measurements from systems  [3]. While the former may not always be accurate enough, the latter may not be feasible for certain tasks. Universal black-box models have already been proposed  [4], but these usually involve large networks that are not suitable for low-latency and real-time applications. In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
In this study, we compare different recurrent artificial neural network architectures for modeling a selection of audio effects. In particular, we considered overdrive, saturator, low-pass filter, equalizer, and compressor. We selected audio effects with different complexities and time dependencies to investigate how these architectural characteristics influence the results when learning directly from raw audio.
@@
For black-box modeling purposes, Recurrent Neural Networks (RNNs) are often preferred among the different types of artificial neural networks. RNNs utilize internal states, which function as memory, making them particularly suitable in scenarios where time dependencies are crucial. Especially when the inputs are limited to a small segment of digital audio samples and the network does not have a large amount of input information to infer, as can happen with feed-forward architectures. Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states. Therefore, dynamic physical systems possess internal states that describe their physical evolution. This similarity to stateful machine learning architectures is often appreciated, as it allows for effective modeling of these systems.
The primary objective of this comparative study is to explore the capabilities of recurrent architectures in learning from raw digital audio and effectively modeling audio effects with different complexities. We aim to determine whether these architectures can learn various types of audio processing and be conditioned by control parameters found in audio effects. Additionally, we investigate whether novel machine learning architectures, which have shown remarkable results in sequence modeling, also perform well in the field of VA.
To address these questions, we apply models based on three different recurrent networks to a variety of audio effects: the popular Long Short-Term Memory (LSTM)  [5], the State-Space models (SSMs)  [6], and the Linear Recurrent Unit (LRU)  [7]. All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  [8] and employ the Gated Linear Unit (GLU)  [9]. For a fair comparison, we ensure that different models have similar numbers of layers and trainable parameters, as well as utilize the same amount of input audio samples.
Moreover, we design these architectures with a limited computational complexity, measured in the number of floating points operations per output audio sample, which is $1.500$ FLOPs per sample or $72$ MFLOPs per second at $48$ KHz, ensuring theoretical real-time computational capabilities on consumer-grade devices. The designed architectures also present limited input-output audio latency, which we have limited to a maximum of $1.33$ ms.
In this study, we consider Convolutional Neural Networks (CNNs) [10] only when integrated with RNNs  [11]. However, when working with small input vectors and considering long-term temporal dependencies, an RNN-based structure with the ability to encode the signal’s history using the states is a particularly effective feature. Therefore, in this study, our focus is on models based on recurrent layers due to their effectiveness in capturing temporal dependencies, and we do not include convolutional-based networks, which are typically more suited for spatial pattern recognition.
@@ 2 Background
There are two possible approaches for approximating a continuous-time system with a discrete-time model: physics-based and data-driven. Physics-based approaches discretize existing continuous-time models. Data-driven approaches identify a discrete model directly from discrete measurement data.
Physics-based modeling is used when comprehensive knowledge about the system is available. In the case of analog audio effects, this requires knowing the schematics of the circuits and specific electronic components populating the system. In addition, the circuit and components need to be mathematically modeled and solvable. A numerical scheme has to be formulated, and the obtained discrete equations have to be solved iteratively to produce the sound of the digital output. In the case of an implicit ordinary differential equations system, the Newton-Raphson (NR) iterative method is generally utilized to find the solutions. NR converges quadratically, although this is not always guaranteed. As a consequence, physics-based models generally present a high computational cost.
@@
Data-driven modeling is used when information on the inner circuitry of the audio effect is not available. The model is considered as a mapping between input-output data: input-output measurements are taken as a reference to find a suitable mapping that approximates the system response. There are three main components in the procedure leading to a data-driven model: the dataset, the set of candidate modeling techniques, and the validation procedure. The selection of input signals for which output is measured, the configuration of the effect’s control parameters, and the data collection process have a major impact on the quality of the resulting model. Data-driven approaches rely on measurements that are static representations of the system behavior, leading to higher difficulty in accurately emulating its dynamical aspects.
Common examples of data-driven approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations  [12]. In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite. These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques. A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in  [13]. These approaches, for the reason of simplification, model linear and nonlinear behavior separately. Using this strategy, two block types can be identified: linear dynamic blocks and static nonlinearities.
Another data-driven approach for modeling analog circuits involves utilizing artificial neural networks (ANNs)  [14]. ANNs are composed of interconnected units, where each connection can transmit a signal to other artificial neurons. These neurons receive and process signals and can transmit them to connected neurons. A real number represents the signal transmitted through a connection, and the output of each neuron is determined by a nonlinear function applied to the sum of its weighted inputs. Neural networks are trained by optimizing the network’s parameters to minimize the error between the predicted output and the actual target values within a given dataset. Gradient-based techniques, such as backpropagation, are commonly employed to estimate the network’s weights. Also, in this scenario, it can be beneficial to incorporate a specific block that focuses on the ”unknown” behavior while separately utilizing a different block to handle the ”known” aspects that have already been accurately modeled.
2.1 Virtual Analog Modeling
Among physics modeling for analog effects, a popular method is based on the wave-digital principle [15]. Wave digital filters (WDF) are constructed by interconnecting simple discrete-time models of individual electronic components (inductors, capacitors, and resistors). Wave digital filters are digital filters based on physical modeling principles. Wave digital (WD) techniques were extensively investigated [16, 17]. Starting with linear elements, WD was expanded to handle nonlinear elements in [18]. More recent advances allowed the realization of multidimensional simulations with multiple nonlinearities  [19]. In [20], authors showed how its quadratic convergence strongly depends on the free parameters and proposed a novel formulation using WD formalism. Recent VA works using this approach include vacuum tube based-circuits as the preamplifier circuit of an amplifier [21], cascaded amplifiers [22], envelope filter guitar effect  [23], booster guitar pedal  [24], and a ring modulator  [25].
Other approaches consider directly deriving the finite-difference time-domain (FDTD) scheme from the continuous physical differential equations. In  [26], this method is used to model an analog tape machine, while in  [27], authors simulate the sonic effect of the Hammond tonewheel organ.
Port-Hamiltonian approach [28] uses the state-space description of power exchange between energy-storing parts, dissipative parts, and external sources in a circuit. Port Hamiltonian formalism has the advantage of preserving the power balance, a beneficial aspect in nonlinear cases  [29]. A port-Hamiltonian formulation is found in  [30], in which a nonlinear passive peaking EQ mode is modeled.
Physics-based methods are popular techniques in circuit modeling. They are more interpretable and flexible (since parametric) but efficiency in this case is not straightforward. Physics modeling can be computationally expensive since they are usually complex formulas and rely on the implicit scheme. In addition, the physics knowledge of the device and in particular the mathematical description may not be challenging to derive which can lead to introducing approximations in the simulated model.
@@ 2.1 Virtual Analog Modeling
Among data-driven block-oriented models, other popular approaches are the block-oriented Hammerstein model, the Wiener model, and their combinations [12]. In the Wiener model, a linear model is followed by a static nonlinearity, while in the Hammerstein model is the opposite. These models have a convenient representation, and they are created by estimating the model coefficients from real data and with some error minimization techniques. A proposed method to classify whether the nonlinear system should rather be modeled by a Wiener or a Hammerstein model is in [13]. Recent examples are for guitar distortion pedal  [31]  [32]  [33], amplifier  [34]  [35], distortion circuit  [36].
@@
In the last two decades, ANNs have been extensively used for modeling analog audio devices. In particular, traditional ML techniques, initially developed for different purposes, have been investigated for creating models that emulate audio effects taking and processing raw audio samples, such as traditional DSP techniques.
In an early attempt using machine learning for audio effect modeling  [37], a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect. The author experimented with a network having one input layer, two hidden layers with sigmoid neurons, and one output layer with a linear activation function. The hidden and output layers present the same size as the input layer, and two different input sizes were considered in the experiments: $512$ and $60$ samples recorded at $44.1$-kHz sampling rate. The training was performed using the MSE as loss function and feeding the network with overlapped segments, in particular, the hop size was half of the input size. The high-pass filter training was carried out on the frequency domain and led to unsatisfactory results, while the distortion effect was trained in the time domain and demonstrated the ability of the network to learn distortion effects. A spurious frequency was found to occur due to the overlapping process since the transition among frames was not sufficiently smooth. Feedforward networks are the simplest type of network, they do not incorporate mechanisms to learn time dependencies but their ability to encode the history of the signal only depends on the amount of input samples they take as input. This aspect limits their use to tasks not involving significant time dependencies.
@@
Other attempts used convolutional-based networks to model the preamplifier circuit from the Fender Bassman 56F-A vacuum-tube amplifier [38]. The dataset has been built using a circuit simulator and the network consists of a stack of dilated causal convolution and fully connected layers. The first fully connected layer uses the gated activation function. The second layer has a hyperbolic tangent activation. Finally, a linear layer outputs the predicted sample. The error-to-signal ratio is the measure used as loss. The selected number of convolutional layers is $10$ with a filter width of three and the dilation pattern $d_{k}$ = ${1,2,4,...,512}$. This configuration leads to a receptive field of $2046$ samples, corresponding to $46$ ms of latency between the input and corresponding output sample, as it uses a sampling rate of $44.1$-kHz. The model presents the fully connected layers with $8$ and the convolutional layers with $16$ units each. The number of samples used as input is set equal to the receptive field. The dataset was simulated with different values of the gain, which is a variable resistor, and used to condition the model. The conditioning is realized in the convolutional layers, and using the gated activation [39]. The work shows that CNNs accurately emulate the reference device better than a fully connected feedforward network and can respond to user control changes.
@@
Using the same architecture, the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals have been modeled as well  [40]. The datasets in this case are collected directly from the devices. Here, the authors investigated more configurations: considering different dilation patterns, activation functions, such as rectified linear unit (ReLU), tanh, and gated activation, numbers of convolutional layers ($10$, $18$, $24$) and number of convolution channels ($2$, $4$, $8$, $16$, and $32$). In this work, the conditioning is not considered. A model with gated activation, $18$ layers, and $16$ convolution channels resulted in lower ESR and real-time capability.
@@
Similarly to a fully connected feedforward network, convolutional networks’ ability to track the past information in the signal depends on the receptive field; a longer receptive field requires a greater number of layers. To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier [41], which present longer time-variant characteristics than distortion units. TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ($2^{2}$ and $02^{1}$ 0). Increasing the dilation factors increases, in turn, the receptive field and the ability of the networks when dealing with long dependencies while, in turn, not increasing the number of layers. In this case, the TCNs have been applied to $44.1$ kHz sampled audio files. A TCN block consists of a convolutional layer, batch normalization, conditioning using FiLM, and PRelu activation function. Authors compared networks with $4$, $5$, and $10$ TCN blocks, $32$ convolutional channel for each convolutional layer, and $5$, $13$, $15$ as the kernel size. The resulting models present receptive fields of $100$, $300$, and $1000$ ms, respectively. The model with a receptive field of $300$ ms provided the best performances. TCNs provide accurate and efficient implementation with respect to computational complexity. On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
@@
A model, that used a combination of different deep learning architectures, was investigated in  [42]. In this case, the model is divided into three parts: an adaptive front-end, latent space, and synthesis back-end. The front-end performs time-domain convolutions with the raw audio to map it into a latent space. It also generates a residual connection, which facilitates the reconstruction of the waveform by the back-end. The adaptive front-end consists of a convolutional encoder. It contains two CNN layers, one pooling layer, and one residual connection. The first layer has $128$ filters and $64$ channels and is followed by the absolute value as a nonlinear activation function. The second layer has $128$ filters of size $128$. The latent-space DNN contains two layers. The first layer is based on locally connected dense layers of $64$ units and the second layer consists of a fully connected layer of $64$ units. Both of these layers are followed by the softplus function. The DNN modifies the latent representation of the input audio into a new latent representation which is fed into the synthesis back-end. The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer. DNN-SAAF consists of four fully connected dense layers of 128, 64, 64, and 128 hidden units, respectively. All dense layers are followed by the softplus function except the last layer. The SAAF [43] is used as the nonlinearity for the last layer. The model is first trained without the latent-space DNN and DNN-SAAF, and then these modules are incorporated for a second training. The loss function to be minimized is the mean absolute error (MAE). In both training procedures, the input and target audio are sliced into frames of $1024$ samples with a hop size of $64$ samples. Distortion, overdrive, and EQ effects have been modeled using this method, with a sampling rate equal to $16$ kHz.
@@
In  [4] the previously described architecture is further explored and compared with different variations and with the CNN architecture based on  [38]. The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space. In this work, the target devices were Universal Audio Vacuum-Tube Preamplifier 610-B, Universal Audio Transistor-Based Limiter Amplifier 1176LN, and 145 Leslie Speaker Cabinet. The highest score in the listening test was obtained by the variation incorporating Bi-LSTM. The input frame size was $4096$ sampled with a hop size of $2048$ samples. Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators  [44]. Here, the loss function to be minimized is the sum of MAE in the time domain and MSE of the log power magnitude spectra of the target and output. These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown to be capable of accurately modeling different types of effects. On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation. In addition, this architecture lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
@@
The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs. These have been applied to model a vacuum-tube amplifier  [45], where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals. The networks consisted of $5$ or $10$ units, while the networks’ input included a variable number of audio samples. The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps. The number of past samples can differ from the input and output signal, and in this case, the experiments considered $5$, $10$, or $50$ input samples and $5$ or $10$ output samples. The work presents preliminary experiments and although the models resulted in poor performance, it demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs. Larger networks generated lower prediction errors, and the overdrive case resulted in more challenges in modeling accurately.
@@
RNNs, in particular LSTMs, have been also compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals  [46]. The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit. LSTM and GRU are compared as RNN layers varying the number of units ($32$, $64$, and $96$). The input signal is extended to include an additional value representing the user control’s parameter. The results show that RNNs achieve similar accuracy to the CNN model while requiring significantly less computational cost and latency since requiring only the current input sample to predict the current output. LSTM performed better than GRU networks, and the error decreased when increasing the number of units. In addition, listening tests showed LSTM networks with $32$ units performing similarly to TCN with $300$ ms of latency  [41].
@@
Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in  [47], which is also compared with a WDF model of the same circuit. ML implementation resulted in better computational performance than the WDF model but sounded very similar although the ML implementation has slightly damped high frequencies. The architecture consists of a GRU layer of $8$ units and one linear fully connected layer with one unit as the output layer. Also, here, a single variable control parameter of the audio effect is added as an extra input to condition the inference of the networks.
@@
A delay-based effect is investigated in  [48], where a tape-based delay is modeled using GRU networks using a $44.1$kHz sample rate. Here, the delay trajectory is analyzed and extracted using impulse train signals. This trajectory is used to demodulate the signal before training the models or used to directly guide a delay-line built in a differentiable way and based on  [1] work. The second strategy produced smaller prediction errors. The authors include the modeling of the tape’s noise and the trajectory using diffusion techniques to approximate the probabilistic distribution. The hidden size of the GRU layer was set to $64$, and attempts with larger sizes led to no improvements.
@@
RNNs rely on their internal memory mechanism to encode the history of the signal and system and allow the design of models to take a small number of input samples. In addition, they are compatible with real-time applications. On the other hand, they can still suffer from remembering significant long-term dependencies. Indeed, LSTMs alone have not been successful for LFO delay-based effects, such as flanger and phaser. In  [49] authors found it beneficial to use as extra input the LFO signal representing the frequency response of a time-varying system over time. The LFO signal is first extracted by analyzing the recording at $44.1$kHz and then used to help the LSTM during the learning process. The extra input has shown to be beneficial to learning the LFO-based effect and can be freely controlled after model training, although in the study is considered only static parameter values and not LFO signal changing depth and width over time. The error-to-signal ratio (ESR) is used as a loss function and the model is an LSTM layer followed by a linear fully connected layer with one unit. The size of the LSTM layer is $16$ or $32$ or $48$.

### v5

# Comparative Study of State-based Neural Networks for Virtual Analog Audio Effects Modeling

@@ Abstract
Analog electronic circuits are at the core of an important category of musical devices, which includes a broad range of sound synthesizers and audio effects. The nonlinear features of their passive and active electronic components give analog musical devices a distinctive timbre and sound quality, making them highly desirable. The development of software that simulates analog musical devices, known as virtual analog modeling, is a significant sub-field in audio signal processing. Artificial neural networks are a promising technique for virtual analog modeling. They have rapidly gained popularity for the emulation of analog audio effects, particularly recurrent networks. While neural approaches have been successful in accurately modeling distortion circuits, they require architectural improvements that account for parameter conditioning and low-latency response. Although hybrid solutions can offer advantages, black-box approaches can still be advantageous in some contexts. In this article, we explore the application of recent machine learning advancements for virtual analog modeling. In particular, we compare State-Space models and Linear Recurrent Units against the more common Long Short-Term Memory networks. These have shown promising ability in sequence-to-sequence modeling tasks, showing a notable improvement in signal history encoding. Our comparative study uses these black-box neural modeling techniques with a variety of audio effects. We evaluate the performance and limitations of these models using multiple metrics, providing insights for future research and development. Our metrics aim to assess the models’ ability to accurately replicate energy envelopes and frequency contents, with a particular focus on transients in the audio signal. To incorporate control parameters into the models, we employ the Feature-wise Linear Modulation method. Long Short-Term Memory networks exhibit better accuracy in emulating distortions and equalizers, while the State-Space model, followed by Long Short-Term Memory networks when integrated in an encoder-decoder structure, and Linear Recurrent Unit outperforms others in emulating saturation and compression. When considering long time-variant characteristics, the State-Space model demonstrates the greatest capability to track history. The Long Short-Term Memory networks tend to introduce audio artifacts.
@@ 1 Introduction
The emulation of analog musical devices has become pervasive in digital products due to their distinct sound, appealing to producers and musicians. Analog circuits use components like operational amplifiers and diodes, whose nonlinear behaviors contribute to creating a unique sound. Replicating these devices, called Virtual Analog (VA) modeling, is now a crucial area in digital audio signal processing.
There are two possible approaches for emulating analog audio effects: physics-based and data-driven. Physics-based approaches discretize existing continuous-time models. Data-driven approaches identify a discrete model directly from discrete measurement data. Machine learning, as a data-driven method, has become widely used in audio modeling and is often integrated into more traditional digital signal-processing frameworks. Hybrid solutions, incorporating knowledge of the modeling tasks, also exist and can offer advantages, but data-driven approaches can still be appealing in some contexts. While the former may not always be accurate enough, the latter may not be feasible for certain tasks.
@@
Data-driven models, also called Black-box, being applicable to different audio modeling tasks, have already been proposed  (Ramírez et al.,, 2020a), but these usually involve large networks that are not suitable for low-latency and real-time applications. In addition, while machine learning modeling of nearly all types of analog effects has been investigated, they often do not provide a satisfactory response when continuously manipulating conditioning parameters, which is needed to offer users runtime control over the effect’s parameters.
@@
This article presents a comparative study that applies recurrent and state-based neural network architectures to model audio effects with different complexities. Neural networks utilizing internal states and a recurrent formulation are often preferred among the different types of artificial neural networks. The most known type is the Recurrent Neural Networks (RNNs). A Similar formulation is present in the State-Space Networks (SSMs). The internal states function as memory, making them particularly suitable in scenarios where time dependencies are crucial. Especially when the inputs are limited to one or a small segment of digital audio samples. For this reason, in this study, we consider Convolutional Neural Networks (CNNs) (LeCun et al.,, 1998) only when integrated with RNNs  (Simionato & Fasciani,, 2022). The aim is to work with small input vectors while also considering long-term temporal dependencies. Therefore, we focus on state-based structures. Analog audio effects are dynamic physical systems that exhibit time-dependent characteristics: their output depends on previous internal states. Therefore, dynamic physical systems possess internal states that describe their physical evolution, similar to stateful machine learning algorithms.
We designed five architectures based on the popular Long Short-Term Memory (LSTM)  (Hochreiter & Schmidhuber,, 1997), the State-Space models (SSMs)  (Gu et al.,, 2022b; Gu & Dao,, 2023), and the Linear Recurrent Unit (LRU)  (Orvieto et al.,, 2023). All models consider conditioning on control parameters using the Feature-wise Linear Modulation (FiLM) method  (Perez et al.,, 2018) and employ the Gated Linear Unit (GLU)  (Dauphin et al.,, 2017). The architectures have a limited computational complexity, measured in the number of floating points operations per output audio sample, ensuring theoretical real-time computational capabilities on consumer-grade devices. The designed architectures also present limited intrinsic input-output audio samples latency, which we have limited to a maximum of $64$ samples ($1.33$ ms). Specifically, we refer to the intrinsic latency given by the architecture constraints, neglecting the latency given by the external DSP processing. This latency is due to the lag between the time index of the input and output samples.
@@ 2 Background
Physics-based modeling for analog audio effects requires comprehensive knowledge of the circuit schematics and components. These circuits need to be mathematically modeled and solved. A numerical scheme is formulated, and the discrete equations are solved iteratively to generate the digital audio output.
@@
Data-driven modeling is used when the inner circuitry of the audio effect is unknown. It treats the model as a mapping between input-output data: input-output measurements are used to approximate the system response. The procedure involves three main components: the dataset, candidate modeling techniques, and the validation process. The quality of the resulting model heavily depends on the selection of input signals, configuration of control parameters, and data collection. Data-driven approaches rely on measurements that are static representations of the system behavior, leading to greater difficulty in accurately emulating its dynamic aspects.
@@
Common examples of data-driven approaches include the block-oriented Hammerstein model, the Wiener model, and their combinations (Schoukens & Tiels,, 2017). In the Wiener model, a linear dynamic system is followed by a static nonlinearity, whereas in the Hammerstein model, the static nonlinearity precedes the linear dynamic system. These models offer a straightforward representation and are derived by estimating the model coefficients from real data using error-minimization techniques. A proposed method to classify whether a Wiener or a Hammerstein model should rather model the nonlinear system is in (Primavera et al.,, 2018). Recent examples are for guitar distortion pedal  (Eichas et al.,, 2015)  (Köper & Holters,, 2020)  (Darabundit et al.,, 2019), amplifier  (Jr & Bodmann,, 2016)  (Eichas et al.,, 2017), distortion circuit  (Eichas & Zölzer,, 2016).
A more recent data-driven technique involves utilizing artificial neural networks (ANNs). ANNs consist of interconnected units called neurons, where each connection can transmit a signal. Each neuron processes these signals and can send them to other connected neurons. The transmitted signal is represented by a real number, and the neuron’s output is determined by applying a nonlinear function to the sum of its weighted inputs. Training ANNs involve optimizing the network’s parameters to minimize the error between predicted and actual target values in a dataset. Gradient-based techniques, like backpropagation, are commonly used to optimize the network’s weights.
2.1 Neural Networks for Virtual Analog Modeling
@@
In the last two decades, ANNs have been extensively used for modeling analog audio devices. In particular, traditional ML techniques, initially developed for different purposes, have been investigated to create models that emulate audio effects when taking and processing raw audio samples, such as traditional DSP techniques.
In an early attempt using machine learning for audio effect modeling  (Mendoza,, 2005), a multilayer feedforward network was utilized to learn a high-pass filter and the Ibanez Tube-Screamer distortion effect. The high-pass filter training led to unsatisfactory results, while the distortion effect demonstrated the ability of the network to learn distortion effects.
@@
Other attempts used convolutional-based networks to model the preamplifier circuit from the Fender Bassman 56F-A vacuum-tube amplifier  (Damskägg et al.,, 2019a). The network consists of a stack of dilated causal convolution and fully connected layers. The dataset was simulated with different values of the gain, which is a variable resistor, and used to condition the model. The conditioning is realized in the convolutional layers and using the gated activation  (Oord et al.,, 2016). The work shows that CNNs accurately emulate the reference device better than a fully connected feedforward network and can respond to user control changes.
@@
Using the same architecture, the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals have been modeled as well  (Damskägg et al.,, 2019b). Here, the authors investigated more configurations, considering different dilation patterns and activation functions. In this work, the conditioning is not considered.
@@
Similarly to a fully connected feedforward network, in which networks’ ability to track the past information depends on the amount of input received, the convolutional network performance depends on the receptive field; a longer receptive field requires a greater number of layers. To overcome this problem, Temporal Convolutional networks (TCN) have been proposed for optical dynamic range compressors such as Teletronix LA-2A Leveling Amplifier (Steinmetz et al.,, 2022), which present longer time-variant characteristics than distortion units. TCNs’ strength consists of dilated convolutional networks with rapidly growing dilation factors ($2^{2}$ and $02^{1}$ 0). Increasing the dilation factors increases, in turn, the receptive field and the ability of the networks when dealing with long dependencies while, in turn, not increasing the number of layers. The authors compared TCN networks with presenting receptive fields of $100$, $300$, and $1000$ ms, respectively. The model with a receptive field of $300$ ms provided the best performances. TCNs provide accurate and efficient implementation with respect to computational complexity. On the other hand, convolutional architectures present an intrinsic latency equal to the length of the receptive field, which is detrimental for live audio applications.
@@
A model that combines different deep learning architectures was investigated in  (Ramírez et al.,, 2019). In this case, the model is divided into three parts: an adaptive front-end, latent space, and synthesis back-end. The front-end performs time-domain convolutions with the raw audio to map it into a latent space. The adaptive front-end consists of a convolutional encoder. The latent-space DNN modifies the latent representation of the input audio into a new latent representation, which is fed into the synthesis back-end. The synthesis back-end consists of an unpooling layer, a deep neural network with smooth adaptive activation functions (DNN-SAAF), and a single CNN layer. The SAAF  (Hou et al.,, 2017) is used as the nonlinearity for the last layer. Distortion, overdrive, and EQ effects have been modeled using this method.
@@
In  (Ramírez et al.,, 2020a), the previously described architecture is further explored and compared with different variations and with the CNN architecture based on  (Damskägg et al.,, 2019a). The variations mainly consist of incorporating Bi-LSTMs or CNNs into the latent space. In this work, the target devices were Universal Audio Vacuum-Tube Preamplifier 610-B, Universal Audio Transistor-Based Limiter Amplifier 1176LN, and 145 Leslie Speaker Cabinet. The highest score in the listening test was obtained by the variation incorporating Bi-LSTM. Another variation based on Bi-LSTM and SAAF as latent-space modules is used for plate and spring reverberators  (Ramírez et al.,, 2020b). These architectures represent a general-purpose deep learning framework for modeling audio effects, which has been shown to be capable of accurately modeling different types of effects. On the other hand, the models are large networks with high computational complexity, leading to difficulty when considering efficient implementation. In addition, using the Bi-LSTM layer makes the architecture non-causal, not allowing real-time applications. This architecture also lacks conditioning possibility and focuses on modeling a static representation of the devices (i.e., fixed parameters).
@@
The last group of machine-learning black-box modeling methods we review in this section are those based on RNNs. These have been applied to model a vacuum-tube amplifier  (Covert & Livingston,, 2013), where authors investigated different networks and input sizes in two scenarios: slightly distorted and overdriven signals. The networks’ input included a variable number of audio samples. The network is autoregressive: it generates the output sample at the current time step using current and past input samples, but also output samples generated at previous time steps. The work presents preliminary experiments, and although the models resulted in poor performance, the work demonstrates the feasibility of modeling vacuum-tube amplifiers with RNNs.
@@
RNNs, in particular LSTMs, have also been compared to CNNs when applied to the Ibanez Tube Screamer, the Boss DS-1, and the Electro-Harmonix Big Muff Pi distortion pedals  (Wright et al.,, 2019). The architecture consists of one RNN layer and one linear fully connected layer as an output layer with one unit. LSTM and GRU are compared as RNN layers varying the number of units. The input signal is extended to include an additional value representing the user control’s parameter. The results show that RNNs achieve similar accuracy to the CNN model while requiring significantly less computational cost and latency since requiring only the current input sample to predict the current output. LSTM performed better than GRU networks, and the error decreased when the number of units was increased.
@@
Another RNN-based model utilized for the Klon Centaur guitar pedal circuit is presented in  (Chowdhury,, 2020), which is also compared with a WDF model of the same circuit. ML implementation resulted in better computational performance than the WDF model but sounded very similar, although the ML implementation slightly damped high frequencies. The architecture consists of a GRU layer and one linear fully connected layer with one unit as the output layer. Also, here, a single variable control parameter of the audio effect is added as an extra input to condition the inference of the networks.
@@
A delay-based effect is investigated in  (Mikkonen et al.,, 2023), where a tape-based delay is modeled using GRU networks. Here, the delay trajectory is analyzed and extracted using impulse train signals. This trajectory is used to demodulate the signal before training the models or used to directly guide a delay-line built in a differentiable way and based on  (Engel et al.,, 2020) work.
@@
RNNs rely on their internal memory mechanism to encode the history of the signal, but they can still suffer from remembering significant long-term dependencies. Indeed, LSTMs alone have not been successful for LFO delay-based effects, such as flanger and phaser. In  (Wright & Välimäki,, 2021), authors found it beneficial to use as extra input the LFO signal representing the frequency response of a time-varying system over time. The LFO signal is first extracted by analyzing the recording and then used to help the LSTM during the learning process. The extra input has shown to be beneficial to learning the LFO-based effect and can be freely controlled after model training, although in the study is considered only static parameter values and not LFO signal changing depth and width over time.